Circular No. 7859

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS A. Henden, Universities Space Research Association and U.S. Naval

Observatory (USNO); U. Munari, Padova; and M. Schwartz, Tenagra Observatory, report that an apparent light echo has been discovered developing around V838 Mon on images secured with the USNO Flagstaff Station 1.0m telescope (scale 0′′.68/pixel) and confirmed on images obtained with the 0.81-m Tenagra telescope (0′′.87/pixel): “The light echo is easily visible in the U bandpass and progressively less visible at the longer B-, V -, R-, and I-band wavelengths due to the combined effect of a decreasing surface brightness and increasing output from the central star. The light echo seems to originate from a circumstellar ring structure, detached from the central source. The light echo was not visible on Feb. 12 images and first appeared on Feb. 17 images. Since then, it has linearly increased in angular diameter at a rate of 0′′.54/day, reaching 27′′ by Mar. 23. Assuming the echo is the result of scattering by a spherical distribution of circumstellar dust of the outburst that started around Jan. 1, the expansion rate suggests a distance to V838 Mon of 700 pc, and first appearance around Feb. 14 suggests a distance of 7800 AU between the inner radius of the circumstellar scattering dust and the central star. If we assume a typical wind velocity of 15 km/s for the material lost by the asymptotic-giant-branch progenitor, the massloss phase that created the circumstellar material stopped 2500 years ago. The central star in V838 Mon appears to be engulfed in a few-arcsec-wide circumstellar cocoon, the possible sign of a recently resumed phase of mass loss. The 45-day delay between onset of the outburst and first detection of the light echo suggests that the sudden brightening of V838 Mon reported on IAUC 7816 should be reaching the circumstellar material shortly and should result in an enhancement of the light echo. V838 Mon seems to be a close analogue of FG Sge, V605 Aql, and V4334 Sgr.”

COMET C/2000 WM_{1} (LINEAR) Visual m_{1} estimates: Feb. 6.30 UT, 3.8 (J. G. de S. Aguiar, Campinas,

Brasil, naked eye); 12.74, 4.4 (D. A. J. Seargent, Cowra, N.S.W., naked eye); 20.52, 5.9 (A. Hale, Cloudcroft, NM, 11×80 binoculars); Mar. 6.83, 7.2 (Y. Nagai, Yamanashi, Japan, 11×80 binoculars); 11.18, 7.6 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 23.84, 8.9 (A. Pearce, Noble Falls, W. Australia, 20×80 binoculars).

2002 March 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green